OpenStudy (anonymous):

find the vertex of the parabola: x = 3y^2 + 6y + 1

OpenStudy (acxbox22):

to find the x value of the vertex, use the formula x=-b/2a plug in the x value into the equation to get the y value then you have your vertex

OpenStudy (anonymous):

can you help me with which ones to plug in

OpenStudy (acxbox22):

sorry but i have to go

OpenStudy (anonymous):

oh alright :/

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

Complete the square: \[\begin{align*}x&=3y^2+6y+1\\ &=3\left(y^2+2y\right)+1\\ &=3\left(y^2+2y+1-1\right)+1\\ &=3\left((y+1)^2-1\right)+1\\ &=3(y+1)^2-3+1\\ &=3(y+1)^2-2\end{align*}\] Now the parabola is in vertex form. The general form would be \(x=a(y-h)^2+k\), where \((h,k)\) is the vertex.

OpenStudy (anonymous):

so the center would be (-1, 2) ?

OpenStudy (anonymous):

vertex* @SithsAndGiggles

OpenStudy (anonymous):

Close, it's (-1,-2).